# Modeling the amplification of epidemic spread by misinformed populations

### v2

@@ Abstract
Understanding how misinformation affects the spread of disease is crucial for public health, especially given recent research indicating that misinformation can increase vaccine hesitancy and discourage vaccine uptake. However, it is difficult to investigate the interaction between misinformation and epidemic outcomes due to the dearth of data-informed holistic epidemic models. Here, we propose an epidemic model that incorporates a large, mobility-informed physical contact network as well as the distribution of misinformed individuals across counties derived from social media data. Our model allows us to simulate and estimate various scenarios to understand the impact of misinformation on epidemic spreading. Using this model, we estimate that misinformation could have led to 47 million additional COVID-19 infections in the U.S. in a worst-case scenario.
Social factors, such as information sharing, play a crucial role in shaping the dynamics and epidemiology of infectious diseases12, 5. For instance, a population’s willingness to adopt public health measures (or lack thereof) largely determines their successes or failures32, 4. A population’s behavioral response to outbreaks can be influenced by mass media, as witnessed during the 2009 H1N1 influenza pandemic38, or by social media and the anti-vaccination movement1, 24, 10, 13.
In particular, misinformation spreading on social networks has been linked to poor compliance with COVID-19 public health guidance43. Greater exposure to unreliable news articles about COVID-19 vaccines has been linked to an increase in vaccine hesitancy and a decrease in vaccination rates at both state and county levels in the United States40, 37. Exposure to online misinformation has also been shown to increase vaccine hesitancy in laboratory experiments31. This is particularly detrimental during vaccination campaigns as clusters of individuals adopting anti-vaccination opinions can make it challenging for a population to reach herd immunity16, 45. Proper management of epidemic crises in the modern age thus requires the understanding of the complex relationship between the spread of (mis)information through online social networks and the spread of disease through physical contact networks (Fig. 1).
@@
While previous agent-based simulations have shown that misinformation may impede the suppression of epidemics in various ways48, 34, 39, 9, there is a growing need to integrate real-world data to strengthen the connections between simulation results and real-world outcomes. Social media can provide valuable data sources for this purpose6, 49.
We address this challenge by proposing a data-informed epidemic model that takes both the distribution of misinformed individuals and a physical mobility network into account. Using this data, we augment the Susceptible-Infected-Recovered (SIR) model to account for a subpopulation of “misinformed” individuals, who do not heed expert public health guidance. We explore how this group can affect the larger, ordinary population using both a mean-field approximation, which assumes all individuals have an equal chance of interacting, and a multi-level agent-based simulation based on two large, data-informed network: a social network where misinformation spreads and a contact network where the disease can propagate. A contact network of approximately 20 million nodes is constructed by leveraging large-scale Twitter data, county-level voting records, and cellphone mobility data. We incorporate theoretically extreme values of the parameter responsible for the epidemic transmission to evaluate best- and worst-case scenarios about the impact of misinformed individuals on the spread of disease and obtain quantitative bounds on the harm caused by misinformation.
@@ Results
Ordinary individuals are considered to be well-informed about public health guidelines, while misinformed individuals, having been exposed to untrustworthy information, are assumed to be less likely to follow recommended behaviors, such as social distancing, mask-wearing, and vaccination, thereby increasing the risk of infection for themselves and others19. We therefore define two distinct parameters $\beta_{O}$ and $\beta_{M}$ for ordinary and misinformed susceptibles, respectively. We model the risky behavior of the misinformed group by setting $\beta_{M}\geq\beta_{O}$. The recovery rate $\gamma$ is identical for both subpopulations. We refer to this as the Susceptible Misinformed Infected Recovered (SMIR) model.
@@
Fig. 3 illustrates the effects of homophily ($0.5\leq\alpha\leq 1$) for different levels of ordinary transmission ($0.1\leq\beta_{O}\leq 0.4$). For less infectious disease (low $\beta_{O}$), increasing homophily significantly harms the misinformed group (Fig. 3B): the infection remains confined within this group. There is no discernible effect on the ordinary population as long as the two groups interact; when they do not ($\alpha=1$), we observe a sharp reduction in infections (Fig. 3A). For $0.160.12<\beta_{O}<$ 0.16, peak infection scenarios coincide with intermediate homophily levels, as indicated by the black dots in Fig. 3C 35. Under these conditions, while increased homophily decreases infections in the general populace, it significantly worsens outcomes for the misinformed group (compare Fig. 3B and C). As $\beta_{O}$ increases further, while nearly the entire misinformed population becomes infected regardless of $\alpha$ (Fig. 3B), high homophily shields the full population (Fig. 3C): ordinary individuals have a lower risk of becoming infected through interactions with misinformed individuals. In summary, homophily offers greater protection to the ordinary group by isolating misinformed communities, which suffer a greater disease burden, exacerbating health disparities 44, 47.
@@
We leverage a large collection of English-language discussions taking place on Twitter about COVID-19 vaccines21. From approximately nine months of this data (Jan. 4–Sep. 30, 2021), we geolocate over 2 million U.S. users who shared almost 26 million tweets and focus on accounts in 341 U.S. counties containing more than 200 Twitter users each. We also infer an account’s political alignment and whether they shared any likely misinformation (see Methods). Twitter is not representative of the U.S. population, and people also access information in other ways, such as traditional media and word of mouth. However, this empirical social network serves as a data-informed proxy for a large, realistic network through which people share information about the disease.
With this data, we build a directed and weighted information diffusion network, in which an edge $(i\rightarrow j,w)$ indicates that $j$ retweeted $i$ $w$ times. There are various ways to model the infodemic18. We simulate the spread of misinformation on this network, as illustrated in Fig. 4A. Accounts that share or reshare posts containing misinformation are considered misinformed. These accounts function as the initial seeds from which misinformation proliferates to the wider network. Not all users actively participate in content sharing, with only about half of U.S. Twitter users, for instance, engaging in this activity36. But even without active sharing, exposure to misinformation can still influence individual behavior31.
To account for users who may be misinformed through exposure, we employ a single-step linear threshold opinion-spreading process25. This is a simple way to model complex contagion15, 51, 33, according to which individuals require multiple exposures to misinformation before they become misinformed themselves. Let a linear threshold $\phi$ represent the minimum number of misinformed friends needed for an ordinary node to become misinformed. If the total number of misinformed friends of $i$ is greater than or equal to $\phi$, $i$ is marked as misinformed ($M$). The remaining nodes are marked as ordinary susceptibles ($O$). We can interpret $\phi$ as a measure of “resilience” to misinformation; as it increases, individuals require more exposure to misinformation to be converted to the misinformed group. Note that since we explore the full range of $\phi$ values, the following results are unaffected whether the threshold is defined based on the number of users or the number of retweets.
@@
We generate contact networks for different thresholds ($1\leq\phi\leq 20$) to compare the impact of misinformed subpopulations of different sizes. Given a threshold $\phi$ and the corresponding information network, we aim to construct a physical contact network containing empirically calibrated misinformed subpopulations (Fig. 4B). The process begins by selecting a sample of individuals from each county within the information network. As party affiliation has been identified as a risk factor associated with excess mortality during the COVID-19 pandemic27, county samples are constructed to match the percentage of Republicans and Democrats who voted in the 2020 U.S. presidential election. For each county, we add the sampled nodes to the physical network marked as misinformed ($M$) or ordinary susceptible ($O$), based on their label within the retweet network. Sampling with replacement allows us to select individuals such that the overall proportions of Republicans and Democrats match the voting records. A 10% sample leads to $N\approx 20$ million nodes. A network based on a much smaller sample is illustrated in Fig. 1B. This process captures empirical measurements of the ideological split, relative population size, and quantity of misinformed individuals in each county. We add contact network edges by leveraging cell phone mobility data that provides the probability of an individual traveling within and between counties. See Methods for details.
@@
Disease-spreading dynamics on a contact network are simulated using an agent-based version of SMIR (Fig. 4C). Two critical parameters affecting these dynamics are the density of the contact network, captured by its average degree $\bar{k}$, and the transmission probability $p$ (note that $p\bar{k}=\beta$). Mitigation measures such as social distancing decrease $\bar{k}$, while those such as masking and vaccination decrease $p$. A simple way to model the combined effects of misinformation on these behaviors is to set $\bar{k}=25$, a high value corresponding to the average number of daily contacts prior to the COVID-19 pandemic30, and explore a wide range of transmission probabilities to capture worst- and best-case scenarios. To this end, we model the refusal of any mitigation measures by selecting the maximum value $p_{M}=1$ for misinformed individuals. In contrast, we model the adoption of several mitigation measures by selecting an extremely small value $p_{O}=0.01$ for ordinary individuals. These values illustrate two theoretically extreme scenarios. The former portrays a realistic number of interactions during non-pandemic times, accompanied by high transmission rates due to the absence of preventive measures, such as social distancing, mask-wearing, or vaccinations. The latter demonstrates decreased daily interactions and reduced transmission rates resulting from the implementation of these preventive measures. Using the empirically calibrated contact networks in conjunction with these extreme parameters, the simulation approach allows us to bound the best- and worst-case scenarios in a data-informed manner (see Methods for more information).
@@ Discussion
This work presents large-scale simulations based on an epidemic model that accounts for a subpopulation of misinformed individuals, informed by real-world data. Such an approach answers calls from the literature to anchor models in empirical data6, 49, allowing us to explore realistic potential outcomes. While we know that exposure to online health misinformation is associated with risky behaviors such as vaccine hesitancy and refusal37, our model further links these behaviors by a misinformed population to an accelerated spread of disease.
@@
Agent-based simulations of the SMIR model let us study the epidemic on empirically calibrated contact networks, which provide more realistic scenarios compared to the mean-field approach. The model estimates that, in the worst-case scenario, a heavily misinformed population can amplify the peak of the infection by a factor of six and accelerate it by two weeks. This would result in an additional 14% of the population becoming infected — nearly 47 million Americans based on recent U.S. Census data50. The corresponding price tag of vaccine misinformation would be over \$143B, using estimated health care costs associated with COVID-19 in the U.S.3.
These results provide an alarming quantitative bound on the harm of exposure to online vaccine misinformation. They should provide public health authorities as well as social media platforms with heightened motivation to curb vaccine misinformation, despite the difficulties posed by social media design11.
We acknowledge certain limitations in our approach. While individual beliefs and behaviors may vary over time, our model simplifies the scenario by dichotomizing individuals into misinformed and ordinary subpopulations and assuming constant transmission rates. We also assume uniform resilience to misinformation for all individuals during the information diffusion process, although this attribute likely differs across individuals. Future directions could involve more sophisticated models to account for these heterogeneities. For instance, cognitive models of misinformation acceptance7 could be incorporated into the simulation with misinformation exposure data collected from social media. Such integration would enable the transition of individuals from ordinary to misinformed susceptible states throughout the simulation, allowing for a simultaneous examination of opinion and disease dynamics. Some theoretical models have already explored similar approaches and obtained results that align with our findings34, 48. Additional extensions of the model could account for a feedback loop whereby witnessing local infections could drive changes in behaviors equivalent to the transition of individuals out of the misinformed population.
@@ Methods
$\left\{\begin{array}[]{@{}l@{}}\frac{dS_{O}}{dt}=-\beta_{O}S_{O}(I_{O}+I_{M}),\;\;\frac{dI_{O}}{dt}=\beta_{O}S_{O}(I_{O}+I_{M})-\gamma I_{O},\;\;\frac{dR_{O}}{dt}=\gamma I_{O}\\ \frac{dS_{M}}{dt}=-\beta_{M}S_{M}(I_{O}+I_{M}),\;\;\frac{dI_{M}}{dt}=\beta_{M}S_{M}(I_{O}+I_{M})-\gamma I_{M},\;\;\frac{dR_{M}}{dt}=\gamma I_{M}.\end{array}\right.$ (1)
@@
$\left\{\begin{array}[]{@{}l@{}}\frac{dS_{O}}{dt}=-2\beta_{O}S_{O}(I_{O}\alpha+I_{M}(1-\alpha)),\;\;\frac{dI_{O}}{dt}=2\beta_{O}S_{O}(I_{O}\alpha+I_{M}(1-\alpha))-\gamma I_{O},\;\;\frac{dR_{O}}{dt}=\gamma I_{O}\\ \frac{dS_{M}}{dt}=-2\beta_{M}S_{M}(I_{O}(1-\alpha)+I_{M}\alpha),\;\;\frac{dI_{M}}{dt}=2\beta_{M}S_{M}(I_{O}(1-\alpha)+I_{M}\alpha)-\gamma I_{M},\;\;\frac{dR_{M}}{dt}=\gamma I_{M}.\end{array}\right.$ (2)
@@
Twitter posts in the CoVaxxy dataset21 were collected in real time via the stream/filter endpoint of the Twitter Application Programming Interface (API). To capture the online discourse surrounding COVID-19 vaccines in English, a comprehensive set of English-language keywords was carefully curated. Beginning with the initial seeds of “covid” and “vaccine,” a snowball sampling technique22 was used to identify co-occurring relevant keywords in December 202021. The resulting list contained almost 80 keywords, available online20. To confirm the relevance of the collected tweets to the topic of vaccines, we examined the coverage obtained by incrementally adding keywords, starting with the most common ones. Over 90% of the tweets in 2021 contained at least one of the three most common keywords: “vaccine,” “vaccination,” or “vaccinate.” To infer the location of accounts, we used the Carmen Python library23 that leverages self-reported location metadata within user profiles (embedded in tweets). As an account’s location may change over time (captured across multiple tweets), we utilize the most recent location. We geolocate 2,047,800 users residing in all 50 U.S. states, who shared a total of 25,806,856 tweets by mapping self-reported locations to U.S. counties. The information network is constructed from accounts in 341 counties that contain more than 200 Twitter users each. Political alignment is estimated using a third-party list of annotated news sources42, 41. It is averaged across all the sources shared by each account. Nodes with an estimated alignment greater (smaller) than zero are considered Republican (Democrat). We infer the political alignment of some additional accounts, who did not share links to news sources, using a label-propagation algorithm17 on the retweet network. If all of a node’s neighbors have political alignment scores, its score is estimated using the weighted average of its neighbors, with weights based on retweets. The process is iterated until each node without a score has at least one neighbor without a score. Misinformation is defined at the source level. Tweets containing links to articles from a list of low-credibility sources compiled by NewsGuard (score below 60) are labeled as spreading misinformation. This approach is common practice and has been validated in the literature37, 26, 8, 29, 46.
@@
To construct edges in our contact network, we utilize SafeGraph cell-phone mobility data52, which contains information on the number of people residing in over 200K Census-Block-Groups (CBG) who visited 4.3M Points-of-Interest (POI) in the United States. This data has been widely employed to study human mobility patterns during the COVID pandemic. We used the average daily number of individuals moving during 2019, as a reference for business-as-usual mobility, and aggregated all CBGs and POIs at the county level. This aggregation results in a county-by-county matrix $L$, where each element $L_{xy}$ represents the average daily number of individuals in county $x$ moving to county $y$ or vice versa. We then normalized $L_{xy}$ to obtain the average probability of individuals in counties $x$ and $y$ coming into contact, and multiplied by the total number of edges to obtain the expected number of connections between individuals in counties $x$ and $y$: $E_{xy}=\frac{L_{xy}}{\sum_{x^{\prime},y^{\prime}}L_{x^{\prime}y^{\prime}}}\frac{\bar{k}N}{2}$ where the sum is over all county pairs and $\frac{\bar{k}N}{2}$ is the total number of edges. Next, we create a physical contact network with $N$ nodes by following a procedure akin to a stochastic block model28 used to generate networks with localized communities. For each pair of distinct locations $x$ and $y$, we draw $E_{xy}$ edges between random pairs of nodes in $x$ and $y$. Additionally, we draw $E_{xx}$ edges among random pairs of individuals within the same location $x$, representing homogeneous mixing within each county. At the end of the process, the network has the target average degree $\bar{k}$. We use $\bar{k}=25$ and show how this parameter affects the infections in Supplementary Information.
@@
Agent-based SMIR simulations are initiated by randomly selecting 100 misinformed nodes and designating them as infected. The disease spreading dynamics are then simulated for 100 steps, which correspond to days. To align with COVID-19 dynamics, we utilize the CDC’s recommended quarantine period of 5 days as our recovery period14 ($\gamma=0.2$). Each simulation is repeated ten times, and the average outcome is reported.

### v3

@@ Abstract
Understanding how misinformation affects the spread of disease is crucial for public health, especially given recent research indicating that misinformation can increase vaccine hesitancy and discourage vaccine uptake. However, it is difficult to investigate the interaction between misinformation and epidemic outcomes due to the dearth of data-informed holistic epidemic models. Here, we employ an epidemic model that incorporates a large, mobility-informed physical contact network as well as the distribution of misinformed individuals across counties derived from social media data. The model allows us to simulate and estimate various scenarios to understand the impact of misinformation on epidemic spreading. Using this model, we present a worst-case scenario in which a heavily misinformed population would result in an additional 14% of the U.S. population becoming infected over the course of the COVID-19 epidemic, compared to a best-case scenario.
Social factors, such as information sharing, play a crucial role in shaping the dynamics and epidemiology of infectious diseases12, 5. For instance, a population’s willingness to adopt public health measures (or lack thereof) largely determines their successes or failures34, 4. A population’s behavioral response to outbreaks can be influenced by mass media, as witnessed during the 2009 H1N1 influenza pandemic40, or by social media and the anti-vaccination movement1, 26, 10, 13.
A great deal of work has explored how to model the influence of human behaviour on the spread of infectious diseases53, 25. Here we focus on risky behaviors affecting disease transmission that are associated with misinformed individuals. Misinformation spreading on social networks has been linked to poor compliance with COVID-19 public health guidance45. Greater exposure to unreliable news articles about COVID-19 vaccines has been linked to an increase in vaccine hesitancy and a decrease in vaccination rates at both state and county levels in the United States42, 39. Exposure to online misinformation has also been shown to increase vaccine hesitancy in laboratory experiments33. This is particularly detrimental during vaccination campaigns as clusters of individuals adopting anti-vaccination opinions can make it challenging for a population to reach herd immunity17, 47. Proper management of epidemic crises in the modern age thus requires the understanding of the complex relationship between the spread of (mis)information through online social networks and the spread of disease through physical contact networks (Fig. 1).
@@
While previous agent-based simulations have shown that misinformation may impede the suppression of epidemics in various ways50, 36, 41, 9, there is a growing need to integrate real-world data to strengthen the connections between simulation results and real-world outcomes. Social media can provide valuable data sources for this purpose6, 51.
We address this challenge by proposing a data-informed epidemic model that takes both the distribution of misinformed individuals and a physical mobility network into account. Using this data, we augment the Susceptible-Infected-Recovered (SIR) model to account for a subpopulation of “misinformed” individuals. We explore how this group can affect the larger, ordinary population using both a mean-field approximation, which assumes all individuals have an equal chance of interacting, and a multi-level agent-based simulation based on two large, data-informed network: a social network where misinformation spreads and a contact network where the disease can propagate. A contact network of approximately 20 million nodes is constructed by leveraging large-scale Twitter data, county-level voting records, and cellphone mobility data. We incorporate theoretically extreme values of the parameter responsible for the epidemic transmission to evaluate best- and worst-case scenarios about the impact of misinformed individuals on the spread of disease and obtain quantitative bounds on the harm caused by misinformation.
@@ Results
Ordinary individuals are considered to be well-informed about public health guidelines, while misinformed individuals, having been exposed to untrustworthy information, are assumed to be less likely to follow recommended behaviors, such as social distancing, mask-wearing, and vaccination, thereby increasing the risk of infection for themselves and others20. We therefore define two distinct parameters $\beta_{O}$ and $\beta_{M}$ for ordinary and misinformed susceptibles, respectively. We model the risky behavior of the misinformed group by setting $\beta_{M}\geq\beta_{O}$. The recovery rate $\gamma$ is identical for both subpopulations. We refer to this as the Susceptible Misinformed Infected Recovered (SMIR) model.
@@
Fig. 3 illustrates the effects of homophily ($0.5\leq\alpha\leq 1$) for different levels of ordinary transmission ($0.1\leq\beta_{O}\leq 0.4$). For less infectious disease (low $\beta_{O}$), increasing homophily significantly harms the misinformed group (Fig. 3B): the infection remains confined within this group. There is no discernible effect on the ordinary population as long as the two groups interact; when they do not ($\alpha=1$), we observe a sharp reduction in infections (Fig. 3A). For $0.160.12<\beta_{O}<$ 0.16, peak infection scenarios coincide with intermediate homophily levels, as indicated by the black dots in Fig. 3C 37. Under these conditions, while increased homophily decreases infections in the general populace, it significantly worsens outcomes for the misinformed group (compare Fig. 3B and C). As $\beta_{O}$ increases further, while nearly the entire misinformed population becomes infected regardless of $\alpha$ (Fig. 3B), high homophily shields the full population (Fig. 3C): ordinary individuals have a lower risk of becoming infected through interactions with misinformed individuals. In summary, homophily offers greater protection to the ordinary group by isolating misinformed communities, which suffer a greater disease burden, exacerbating health disparities 46, 49.
@@
We leverage a large collection of English-language discussions taking place on Twitter about COVID-19 vaccines22. From approximately nine months of this data (Jan. 4–Sep. 30, 2021), we geolocate over 2 million U.S. users who shared almost 26 million tweets and focus on accounts in 341 U.S. counties containing more than 200 Twitter users each. We also infer an account’s political alignment and whether they shared any likely misinformation (see Methods). Twitter is not representative of the U.S. population, and people also access information in other ways, such as traditional media and word of mouth. However, this empirical social network serves as a data-informed proxy for a large, realistic network through which people share information about the disease.
With this data, we build a directed and weighted information diffusion network, in which an edge $(i\rightarrow j,w)$ indicates that $j$ retweeted $i$ $w$ times. There are various ways to model the infodemic19. We simulate the spread of misinformation on this network, as illustrated in Fig. 4A. Accounts that share or reshare posts containing misinformation are considered misinformed. These accounts function as the initial seeds from which misinformation proliferates to the wider network. Not all users actively participate in content sharing, with only about half of U.S. Twitter users, for instance, engaging in this activity38. But even without active sharing, exposure to misinformation can still influence individual behavior33.
To account for users who may be misinformed through exposure, we employ a single-step linear threshold opinion-spreading process27. While many social influence models have been proposed14, this is a simple way to capture complex contagion, according to which individuals require multiple exposures to misinformation before they become misinformed themselves16, 54, 35. Let a linear threshold $\phi$ represent the minimum number of misinformed friends needed for an ordinary node to become misinformed. If the total number of misinformed friends of $i$ is greater than or equal to $\phi$, $i$ is marked as misinformed ($M$). The remaining nodes are marked as ordinary susceptibles ($O$). We can interpret $\phi$ as a measure of “resilience” to misinformation; as it increases, individuals require more exposure to misinformation to be converted to the misinformed group. Note that since we explore the full range of $\phi$ values, the following results are unaffected whether the threshold is defined based on the number of users or the number of retweets.
@@
We generate contact networks for different thresholds ($1\leq\phi\leq 20$) to compare the impact of misinformed subpopulations of different sizes. Given a threshold $\phi$ and the corresponding information network, we aim to construct a physical contact network containing empirically calibrated misinformed subpopulations (Fig. 4B). The process begins by selecting a sample of individuals from each county within the information network. As party affiliation has been identified as a risk factor associated with excess mortality during the COVID-19 pandemic29, county samples are constructed to match the percentage of Republicans and Democrats who voted in the 2020 U.S. presidential election. For each county, we add the sampled nodes to the physical network marked as misinformed ($M$) or ordinary susceptible ($O$), based on their label within the retweet network. Sampling with replacement allows us to select individuals such that the overall proportions of Republicans and Democrats match the voting records. A 10% sample leads to $N\approx 20$ million nodes. A network based on a much smaller sample is illustrated in Fig. 1B. This process captures empirical measurements of the ideological split, relative population size, and quantity of misinformed individuals in each county. We add contact network edges by leveraging cell phone mobility data that provides the probability of an individual traveling within and between counties. See Methods for details.
@@
Disease-spreading dynamics on a contact network are simulated using an agent-based version of SMIR (Fig. 4C). Two critical parameters affecting these dynamics are the density of the contact network, captured by its average degree $\bar{k}$, and the transmission probability $p$ (note that $p\bar{k}=\beta$). Mitigation measures such as social distancing decrease $\bar{k}$, while those such as masking and vaccination decrease $p$. A simple way to model the combined effects of misinformation on these behaviors is to set $\bar{k}=25$, a high value corresponding to the average number of daily contacts prior to the COVID-19 pandemic32, and explore a wide range of transmission probabilities to capture worst- and best-case scenarios. To this end, we model the refusal of any mitigation measures by selecting the maximum value $p_{M}=1$ for misinformed individuals. In contrast, we model the adoption of several mitigation measures by selecting an extremely small value $p_{O}=0.01$ for ordinary individuals. These values illustrate two theoretically extreme scenarios. The former portrays a realistic number of interactions during non-pandemic times, accompanied by high transmission rates due to the absence of preventive measures, such as social distancing, mask-wearing, or vaccinations. The latter demonstrates decreased daily interactions and reduced transmission rates resulting from the implementation of these preventive measures. Using the empirically calibrated contact networks in conjunction with these extreme parameters, the simulation approach allows us to bound the best- and worst-case scenarios in a data-informed manner (see Methods for more information).
@@ Discussion
This work presents large-scale simulations based on an epidemic model that accounts for a subpopulation of misinformed individuals, informed by real-world data. Such an approach answers calls from the literature to anchor models in empirical data6, 51, allowing us to explore realistic potential outcomes. While we know that exposure to online health misinformation is associated with risky behaviors such as vaccine hesitancy and refusal39, our model further links these behaviors by a misinformed population to an accelerated spread of disease.
@@
Agent-based simulations of the SMIR model let us study the epidemic on empirically calibrated contact networks, which provide more realistic scenarios compared to the mean-field approach. The model estimates that, in the worst-case scenario, a heavily misinformed population can amplify the peak of the infection by a factor of six and accelerate it by two weeks. This would result in an additional 14% of the population becoming infected — nearly 47 million Americans based on recent U.S. Census data52. The corresponding price tag of vaccine misinformation would be over \$143B, using estimated health care costs associated with COVID-19 in the U.S.3.
While these figures are based on extreme scenarios, they represent an alarming bound on the harm of exposure to online vaccine misinformation. They should provide public health authorities as well as social media platforms with heightened motivation to curb vaccine misinformation, despite the difficulties posed by social media design11.
We acknowledge certain limitations in our approach. While individual beliefs and behaviors may vary over time, our model simplifies the scenario by dichotomizing individuals into misinformed and ordinary subpopulations and assuming constant transmission rates. We also assume uniform resilience to misinformation for all individuals during the information diffusion process, although this attribute likely differs across individuals. Future directions could involve more sophisticated models to account for these heterogeneities. For instance, cognitive models of misinformation acceptance7 could be incorporated into the simulation with misinformation exposure data collected from social media. Such integration would enable the transition of individuals from ordinary to misinformed susceptible states throughout the simulation, allowing for a simultaneous examination of opinion and disease dynamics. Some theoretical models have already explored similar approaches and obtained results that align with our findings36, 50.
Future work should attempt to analyze the potential effects of misinformation in more realistic scenarios, where values for the key parameters $p_{M}$ and $p_{O}$ could be obtained by calibrating the model with empirical epidemic data from particular regions and time periods. Additional extensions of the model could account for a feedback loop whereby witnessing local infections could drive changes in behaviors equivalent to the transition of individuals out of the misinformed population.
@@ Methods
$\left\{\begin{array}[]{@{}l@{}}\frac{dS_{O}}{dt}=-\beta_{O}S_{O}(I_{O}+I_{M}),\;\;\frac{dI_{O}}{dt}=\beta_{O}S_{O}(I_{O}+I_{M})-\gamma I_{O},\;\;\frac{dR_{O}}{dt}=\gamma I_{O}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \frac{dS_{M}}{dt}=-\beta_{M}S_{M}(I_{O}+I_{M}),\;\;\frac{dI_{M}}{dt}=\beta_{M}S_{M}(I_{O}+I_{M})-\gamma I_{M},\;\;\frac{dR_{M}}{dt}=\gamma I_{M}.\end{array}\right.$ (1)
@@
[…]
@@
Twitter posts in the CoVaxxy dataset22 were collected in real time via the stream/filter endpoint of the Twitter Application Programming Interface (API). To capture the online discourse surrounding COVID-19 vaccines in English, a comprehensive set of English-language keywords was carefully curated. Beginning with the initial seeds of “covid” and “vaccine,” a snowball sampling technique23 was used to identify co-occurring relevant keywords in December 202022. The resulting list contained almost 80 keywords, available online21. To confirm the relevance of the collected tweets to the topic of vaccines, we examined the coverage obtained by incrementally adding keywords, starting with the most common ones. Over 90% of the tweets in 2021 contained at least one of the three most common keywords: “vaccine,” “vaccination,” or “vaccinate.” To infer the location of accounts, we used the Carmen Python library24 that leverages self-reported location metadata within user profiles (embedded in tweets). As an account’s location may change over time (captured across multiple tweets), we utilize the most recent location. We geolocate 2,047,800 users residing in all 50 U.S. states, who shared a total of 25,806,856 tweets by mapping self-reported locations to U.S. counties. The information network is constructed from accounts in 341 counties that contain more than 200 Twitter users each. Political alignment is estimated using a third-party list of annotated news sources44, 43. It is averaged across all the sources shared by each account. Nodes with an estimated alignment greater (smaller) than zero are considered Republican (Democrat). We infer the political alignment of some additional accounts, who did not share links to news sources, using a label-propagation algorithm18 on the retweet network. If all of a node’s neighbors have political alignment scores, its score is estimated using the weighted average of its neighbors, with weights based on retweets. The process is iterated until each node without a score has at least one neighbor without a score. Misinformation is defined at the source level. Tweets containing links to articles from a list of low-credibility sources compiled by NewsGuard (score below 60) are labeled as spreading misinformation. This approach is common practice and has been validated in the literature39, 28, 8, 31, 48.
@@
To construct edges in our contact network, we utilize SafeGraph cell-phone mobility data55, which contains information on the number of people residing in over 200K Census-Block-Groups (CBG) who visited 4.3M Points-of-Interest (POI) in the United States. This data has been widely employed to study human mobility patterns during the COVID pandemic. We used the average daily number of individuals moving during 2019, as a reference for business-as-usual mobility, and aggregated all CBGs and POIs at the county level. This aggregation results in a county-by-county matrix $L$, where each element $L_{xy}$ represents the average daily number of individuals in county $x$ moving to county $y$ or vice versa. We then normalized $L_{xy}$ to obtain the average probability of individuals in counties $x$ and $y$ coming into contact, and multiplied by the total number of edges to obtain the expected number of connections between individuals in counties $x$ and $y$: $E_{xy}=\frac{L_{xy}}{\sum_{x^{\prime},y^{\prime}}L_{x^{\prime}y^{\prime}}}\frac{\bar{k}N}{2}$ where the sum is over all county pairs and $\frac{\bar{k}N}{2}$ is the total number of edges. Next, we create a physical contact network with $N$ nodes by following a procedure akin to a stochastic block model30 used to generate networks with localized communities. For each pair of distinct locations $x$ and $y$, we draw $E_{xy}$ edges between random pairs of nodes in $x$ and $y$. Additionally, we draw $E_{xx}$ edges among random pairs of individuals within the same location $x$, representing homogeneous mixing within each county. At the end of the process, the network has the target average degree $\bar{k}$. We use $\bar{k}=25$ and show how this parameter affects the infections in Supplementary Information.
@@
Agent-based SMIR simulations are initiated by randomly selecting 100 misinformed nodes and designating them as infected. The disease spreading dynamics are then simulated for 100 steps, which correspond to days. To align with COVID-19 dynamics, we utilize the CDC’s recommended quarantine period of 5 days as our recovery period15 ($\gamma=0.2$). Each simulation is repeated ten times, and the average outcome is reported.